Foreverangel:

Can someone check if my answer is correct or not please? Answer: 30 -17 51 6 2 -24 18 -3 -15

Foreverangel:

1 attachment
Serenity1jacksonsl:

no it is wrong you did not put it correctly

surjithayer:

\[\left[\begin{matrix}8 & -7 \\ -2 & 4\\0 &3\end{matrix}\right]\left[\begin{matrix}9 & -3&2 \\ 6 & -1&-5\end{matrix}\right]\] \[=\left[\begin{matrix}8*9+(-7)*6 & 8*(-3)+(-7)*(-1)&8*2+(-7)* (-5) \\ -2*9+4*6 & (-2)*(-3)+4*(-1)&-2*2+4*(-5)\\0*9+3*6&0*(-3)+3*(-1)&0*2+3*(-5)\end{matrix}\right]\] \[=\left[\begin{matrix}72-42 & -24+7&16+35 \\ -18+24 & 6-4&-4-20\\0+18&0-3&0-15\end{matrix}\right]\] \[=\left[\begin{matrix}30 & -17&51 \\ 6 & 2&-24\\18&-3&-15\end{matrix}\right]\]

surjithayer:

PQ is correct.

surjithayer:

QP will be differnt.

Foreverangel:

@surjithayer wrote:
QP will be differnt.
Sorry I forgot to put that it was looking for PQ but thank you for checking!